OpenStudy (anonymous):

2. A 4.0 kg ball moving to the right at 1.2 m/s collides inelastically with a 1.0 kg ball at rest. What will the velocity of the combined balls be after the collision? 3. An object is lowered into a deep hole in the ground. How does the potential energy of the object change? 4. A real spring does not oscillate forever. Instead, it eventually comes to a stop. Does this violate the Law of Conservation of Energy? Explain why or why not. 5. A boy can raise a rock that weighs 100 N by using a lever and applying a force of 20 N. What is the mechanical advantage of the lever?

OpenStudy (anonymous):

6. What principle explains why lifting heavy objects is easier using ramps?

OpenStudy (anonymous):

2.in an inelastic collision,the colliding bodies move together as a single body after collision. use conservation of momentum. 3.potential energy of an object is given by=mgh where h is the height of that object from any reference level (where the potential energy is assumed to be zero). 4.No.it does not violate the law of conservation of energy.the spring eventually stops because the energy of the spring is lost in the form of heat due to air resistance and other factors of our surroundings 5.the lever reduces the amount of mechanical force that boy needs to exert for lifting any block

OpenStudy (anonymous):

5. MA = 100/20 = 5/1